y ln x ln y dx + dy =0


1
Expert's answer
2021-09-01T08:28:01-0400

ylnxlnydx=-dy

lnxdx=-dy/(ylny)

"\intop lnxdx=fg-\int f'g"

f=lnx

g'=1

f'=1/x

g=x

"\int lnxdx=xlnx- \int 1dx=x(lnx-1)+C"

"\int \frac{dy}{ylny}"

u=lny

du/dy=1/y

dy=ydu"\int \frac{1}{u} du= lnu"

"\int \frac{dy}{ylny}=ln(lny)+C"

"x(lnx-1)+C=ln(lny)"

"y=e^{e^{x(lnx-1)+C}}"